i need to find the equation of a line that is perpendicular to the line y=1/4x+6 and contains the points (-8,0)
Answer by scott8148(6628)

y = (1/4)x + 6 is in slope-intercept form (y = mx + b), with 1/4 being the slope

perpendicular lines have slopes that are negative reciprocals, so the slope of the new line is -4

y = -4 x + b

to find the new y-intercept (b), plug in the given point

(0) = -4(-8) + b ___ -32 = b

the new equation is ___ y = -4x - 32